OpenStudy (anonymous):

i love kitties Match each item in Column A to an answer in Column B. 2. (1 pt) What is the distance between the two points on a number line? Column A 1. X = 17, Y = 3 2. P = –14, Q = –6 3. U = –5, V = –13 4. J = 9, K = –11 Column B A. 8 B. 14 C. 20

OpenStudy (anonymous):

@imsleeziboii @TheCalcHater

OpenStudy (thecalchater):

post the pics

OpenStudy (imsleeziboii):

Is there pictures?

OpenStudy (anonymous):

there isnt one

OpenStudy (imsleeziboii):

Wheres the number lines?

OpenStudy (thecalchater):

Do you know the distance formula?

OpenStudy (anonymous):

no

OpenStudy (imsleeziboii):

Is there number lines? we need the points

OpenStudy (anonymous):

no it doesnt have one

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (imsleeziboii):

That doesnt make any sense it asks to match the points

OpenStudy (thecalchater):

C. 20 Dont forget to fan and medal

OpenStudy (imsleeziboii):

Oh i see now

OpenStudy (anonymous):

c20 isnt an answer

OpenStudy (anonymous):

whatis it :'-()

OpenStudy (imsleeziboii):

i gotcha one sec

OpenStudy (imsleeziboii):

1.Difference of 10

OpenStudy (imsleeziboii):

wait idk

OpenStudy (imsleeziboii):

sorry ill tag someone

OpenStudy (imsleeziboii):

@ShadowLegendX

OpenStudy (trojanpoem):

X = 17 Y =3 Distance = |17-3| = 14 Can you continue the rest ?

OpenStudy (trojanpoem):

P = -14 Q = -6 Distane = | -14 - (-6)| = |-14 + 6| = |-8| = 8

OpenStudy (anonymous):

b a a c is this right

OpenStudy (trojanpoem):

U = –5, V = –13 4. J = 9, K = –11 Distance U,V = |-5 +13| = |8| = 8 Distance J,K = | 9 - (-11) | = |9 + 11 | = |20| = 20

OpenStudy (trojanpoem):

so 14 8 8 20